# Non-Linear Matter Power Spectrum Modeling in Interacting Dark Energy Cosmologies

### v1

@@ Abstract
Understanding the behavior of the matter power spectrum on non-linear scales beyond the $\Lambda$CDM model is crucial for accurately predicting the large-scale structure (LSS) of the Universe in non-standard cosmologies. In this work, we present an analysis of the non-linear matter power spectrum within the framework of interacting dark energy-dark matter cosmologies (IDE). We employ N-body simulations and theoretical models to investigate the impact of IDE on these non-linear scales. Beginning with N-body simulations characterized by a fixed parameter space delineated by prior observational research, we adeptly fit the simulated spectra with a simple parametric function, achieving accuracy within 5%. Subsequently, we refine a modified halo model tailored to the IDE cosmology, exhibiting exceptional precision in fitting the simulations down to scales of approximately 1 h/Mpc. To assess the model’s robustness, we conduct a forecast analysis for the Euclid survey, employing our refined model. We find that the coupling parameter $\xi$ will be constrained to $\sigma(\xi)=0.0046$. This marks a significant improvement by an order of magnitude compared to any other current observational tests documented in the literature. These primary findings pave the way for a novel preliminary approach, enabling the utilization of IDE models for observational constraints concerning LSS data on non-linear scales.
@@ I Introduction
The non-linear power spectrum is well modeled for the $\Lambda$CDM model and some of its basic extensions Takahashi et al. (2012); Matsubara (2008); Hand et al. (2017); Vlah et al. (2016); Chen et al. (2021); Baumann et al. (2012); Cabass et al. (2023). However, modeling non-linear regimes in non-standard cosmologies poses a significant challenge. Various methodologies have been employed to address this challenge, including the Effective Field Theory of LSS D’Amico et al. (2021a, b); Ivanov et al. (2020); Colas et al. (2020); D’Amico et al. (2020); Bose et al. (2018), the Halo models Bose et al. (2023); Carrilho et al. (2022); Bose et al. (2020); Gupta et al. (2023); Carrion et al. (2024); Spurio Mancini and Pourtsidou (2022), and emulators trained via N-body simulation data Winther et al. (2019); Fiorini et al. (2023); Mauland et al. (2023); Sáez-Casares et al. (2024); Ruan et al. (2024). These approaches offer promising avenues for improving our understanding of non-linear regimes in cosmology, but further research is needed to explore their capabilities and limitations.
Given the imperative to anticipate model behavior on non-linear scales, it becomes indispensable to employ phenomenological models for the interpretation of current and forthcoming data derived from galaxy clustering and cosmic shear. By doing so, we can better understand the perturbation dynamics at play within these complex systems and extract meaningful insights from observational data.
In this work, our main aim is to build a robust non-linear $P(k)$ model in the presence of a DE-DM interaction. To this end, we carry out a series of N-body gravity simulations and then we develop these perspectives in two ways. First, we capture the ratio $R(k)=P_{\rm NL}(k)/P_{\rm linear}(k)$ up to non-linear scales, and we find reliable analytical modeling for the function $R(k)$ up to scales $k\leq 1$ h/Mpc. In a second modeling context, we modify and propose a new $P(k)$ Halo model, which can fit the simulation data. Finally, we applied our models in a forecast analysis from the perspective of the Euclid survey. We demonstrate that an Euclid-like survey will possess strong sensitivity to constrain DM-DE coupling, rendering it a crucial probe for this particular class of cosmological models.
@@ II Background expansion and linear perturbations
Let us start by examining the fundamental characteristics of IDE models. Throughout our discussion, we will operate on the assumption of a spatially flat Friedmann-Lemaître-Robertson-Walker (FLRW) metric. We can introduce a phenomenological parameterization for the DM-DE interaction within the conservation equations such that the individual stress-energy tensors are no longer conserved, although their sum remains so. [^1] See Nunes et al. (2022a) for a more detailed description and Wang et al. (2024b, 2016) for a general review. Typically, it is assumed that the covariant derivatives of the DM and DE stress-energy tensors evolve according to:
@@
The introduction of the DM-DE coupling further alters the evolution of perturbations. Working in the Newtonian gauge, the coupled system of linear Einstein-Boltzmann equations governing the evolution of dark matter density perturbations, $\delta_{c}$, and velocity divergences, $\theta_{c}$, is expressed as follows Väliviita et al. (2008); Honorez et al. (2010); Gavela et al. (2010):
@@
It is worth noting that the Euler equation remains unaltered compared to $\Lambda$CDM, indicating the absence of a fifth force in this model. Additionally, we will not consider the potential presence of clustering DE in this study. As argued in Dakin et al. (2019), DE perturbations may only have discernible effects on very large scales, necessitating a relativistic treatment of simulations. In the regime of Newtonian N-body simulations that we are focused on modeling, DE clustering does not significantly impact the matter power spectrum Hassani et al. (2019). Moreover, recent analyses Dinda and Banerjee (2023) have demonstrated that observations favor a homogeneous DE component over a clustering one. However, the presence of a homogeneous DE component influences simulations in various ways, including background evolution, growth function calculation, and initial conditions. To provide a more precise quantification within the IDE context, Figure 1 illustrates the relative difference in $P(k,z=0)$ with and without DE clustering effects. In constructing the figure, we adopted IDE-01 and IDE-02 baseline values, which will be defined in the subsequent section. It is evident that on large scales, the observed effect is minimal, amounting to less than 1%. As for scales where non-linear effects are expected to dominate (i.e., $k>0.2h/\text{Mpc}$), the relative difference is approximately 1%. Consequently, we can infer that the perturbations due to DE can be considered negligible. The results were obtained using a customized version of the CLASS code Blas et al. (2011), which integrates the IDE model elaborated upon in this study.
@@
Figure 2: The growth factor is depicted as a function of redshift. The dashed black curve represents the standard model, $\Lambda$CDM, whereas the solid curves correspond to the interacting model utilized, showcasing variations of the $\xi$ parameter.
@@ III.1 Methodology
The cosmic structure formation in a given theory is most accurately modeled with N-body simulations, where the particles in the simulation are incrementally displaced from their initial positions according to the Poisson equation, the expansion rate of the universe and initial conditions. In this work, we will use the pmwd code Li et al. (2022), a differentiable cosmological particle-mesh N-body library.
@@
To test our framework for the IDE model in N-body simulations introduced above, we shall consider three scenarios: the $\Lambda$CDM model from Planck-CMB best-fit values, and, based on the recent LSS analysis Nunes et al. (2022b) in the IDE context, two different input baselines for IDE. Therefore, we will consider the following common fiducial values: $2.110^{9}A_{s}=$ 2.1, $n_{s}=0.96$, and $\Omega_{b}=0.05$, while we will choose for:
@@ III.2 Discussion
In the following discussion, the superscript $s$ indicates quantities evaluated differently from N-body simulation data. Figure 4 illustrates the ratio between $R_{\text{IDE}}^{s}(k)$ and $R_{\text{IDE}}(k)$, parameterized by equation (31). Our analysis reveals that the parameters $A\in[23.23,24.24]$, $B\in[3.77,4.74]$, $C\in[-0.03,-0.04]$, and $D\in[-9.30,-16.35]$ can effectively fit the non-linear spectra within a $<$5% relative difference for the values within the range adopted in IDE-01 and IDE-02 samples. It is noteworthy that any observed noise in Figure 4 on very large scales is solely attributable to the box size of our simulations. Thus, for the regime where non-linear effects are predominant, i.e., $k>0.2$ h/Mpc, we note that our fit is very precise with accuracy $<$1%.
@@ IV The Halo Model for IDE
The leading order moment of the cosmological matter distribution is the non-linear matter power spectrum, $P_{\rm NL}(k)$. We follow the halo model (see Asgari et al. (2023) for a review) based approach of HMCode model Mead et al. (2016), in a target theory of cosmology to modeling $P_{\rm NL}(k)$ for the IDE scenario under consideration in this work.
@@
In the above equation, the power spectrum is formulated as an integral over halo masses, $M$, with the differential halo mass function denoted as $dM$, determined by the halo mass function, $F(M)$. Additionally, this integral incorporates the term $W(k,M)$, representing the normalized Fourier transform of the halo density profile Asgari et al. (2023). On larger spatial scales, the distribution of haloes exhibits non-random patterns, and the displacements between them necessitate the incorporation of a two-halo term into the power spectrum analysis. This term accounts for the interactions between haloes and is akin to the linear-theory power spectrum for the matter distribution, $\Delta_{2\mathrm{H}}^{2}(k)=\Delta_{\mathrm{lin}}^{2}(k)$.
@@
Our analysis indicates that $\beta_{1}\in[3.4,3.6]$, $\beta_{2}\in[0.15,0.28]$, $\beta_{3}\in[0.15,1.31]$, $\beta_{4}\in[3.5,5.1]$ and $\beta_{5}\in[1.09,1.11]$ can fit with good accuracy the entire parameter-space range between the IDE-01 and IDE-02 model baselines. The creation of these parameters played a crucial role in our model’s ability to accurately describe the power spectra of matter on non-linear scales. Figure 5 illustrates the ratio between the modified halo model and simulation data for two scenarios (IDE-01 and IDE-02). Notably, modeling accuracy within a 5% margin extends to values of $k\approx 1$. A relative difference of less than 5% compared to simulations is a promising outcome, suggesting that our model may be robust enough to capture the subtleties of matter structures even in different settings, as tested with IDE-01 and IDE-02.
@@ V Forecast analysis from the Euclid Survey perspective
Parameter $\sigma({\rm Euclid})$
$\Omega_{\rm cdm}$ 0.0013
$n_{s}$ 0.0043
$H_{0}$ 0.55
$\xi$ 0.0046
$\sigma_{\rm NL}$ 0.054
$b_{0}$ 0.0034
$b_{1}$ 0.0060
@@
In what follows in our main results, we use a modified version of the CLASS$+$MontePython code Blas et al. (2011); Audren et al. (2013); Brinckmann and Lesgourgues (2019), adapting the likelihoods euclid_pk and euclid_lensing as well as incorporating all background dynamics, linear perturbations, and the modified halo models in CLASS, as discussed in the previous sections. In our forecast analyses, we utilize a maximum scale of $k=1$ h/Mpc, within which our model demonstrates good predictive capability in fitting the simulations.
@@
Of particular significance is the examination of sensitivity towards the coupling parameter $\xi$. According to the analysis of Euclid forecasts, the coupling parameter can be constrained with a precision of $\sigma(\xi)=0.0046$. This represents a substantial improvement by an order of magnitude compared to any other current observational tests documented in the literature. For example, in a recent study Nunes et al. (2022b), which employed the full-shape galaxy power spectrum from the BOSS DR12 sample in conjunction with CMB data from Planck, a constraint of $\xi>-0.12$ was established. Consequently, the forthcoming data from Euclid has the potential to enhance these constraints by up to two orders of magnitude. Therefore, the accuracy of the constraints generated by the Euclid mission sensitivity will be crucial for the IDE models. The joint accuracy with $\sigma(\Omega_{\rm cdm})$ and $\sigma(H_{0})$, two parameters highly sensitive to IDE dynamics, will also be constrained with remarkable precision. Consequently, Euclid data alone will possess the capability to constrain the dynamics of IDE models, encompassing both background and perturbative levels, with unprecedented accuracy.
@@
There are other noteworthy points to consider. In the forecast analysis from the perspective of Euclid, it is observed that the well-established correlation between $\xi$ and $H_{0}$ persists, indicating a significant negative correlation. This correlation has played a crucial role in arguing that this particular class of models may offer a resolution to the $H_{0}$ tension Zhai et al. (2023); Bernui et al. (2023); Gariazzo et al. (2022). Additionally, it is intriguing to observe that $\xi$ may exhibit a notable new correlation with the galaxy bias parameters $b_{0}$ and $b_{1}$ (see Figure 6). Furthermore, the correlation in the $\xi$-$H_{0}$ plane within the IDE framework implies that the galaxy bias parameters will also exhibit correlations with $H_{0}$. Consequently, this reveals a previously unrecognized cosmological characteristic for this class of models, which will be essential to comprehend for future analyses with real observational data.

### v2

@@ Abstract
Understanding the behavior of the matter power spectrum on non-linear scales beyond the $\Lambda$CDM model is crucial for accurately predicting the large-scale structure (LSS) of the Universe in non-standard cosmologies. In this work, we present an analysis of the non-linear matter power spectrum within the framework of interacting dark energy-dark matter cosmologies (IDE). We employ N-body simulations and theoretical models to investigate the impact of IDE on these non-linear scales. Beginning with N-body simulations characterized by a fixed parameter space delineated by prior observational research, we adeptly fit the simulated spectra with a simple parametric function, achieving accuracy within 5%. Subsequently, we refine a modified halo model tailored to the IDE cosmology, exhibiting exceptional precision in fitting the simulations down to scales of approximately 1 h/Mpc. To assess the model’s robustness, we conduct a forecast analysis for the Euclid survey, employing our refined model. We find that the coupling parameter $\xi$ will be constrained to $\sigma(\xi)=0.0110$. This marks a significant improvement by an order of magnitude compared to any other current observational tests documented in the literature. These primary findings pave the way for a novel preliminary approach, enabling the utilization of IDE models for observational constraints concerning LSS data on non-linear scales.
@@ I Introduction
The non-linear power spectrum is well modeled for the $\Lambda$CDM model and some of its basic extensions Takahashi et al. (2012); Matsubara (2008); Hand et al. (2017); Vlah et al. (2016); Chen et al. (2021); Baumann et al. (2012); Cabass et al. (2023). However, modeling non-linear regimes in non-standard cosmologies poses a significant challenge. Various methodologies have been employed to address this challenge, including the Effective Field Theory of LSS D’Amico et al. (2021a, b); Ivanov et al. (2020); Colas et al. (2020); D’Amico et al. (2024); Bose et al. (2018), the Halo models Bose et al. (2022); Carrilho et al. (2022); Bose et al. (2020); Gupta et al. (2023); Carrion et al. (2024); Mancini Spurio and Pourtsidou (2022), and emulators trained via N-body simulation data Winther et al. (2019); Fiorini et al. (2023); Mauland et al. (2024); Sáez-Casares et al. (2023); Ruan et al. (2024). These approaches offer promising avenues for improving our understanding of non-linear regimes in cosmology, but further research is needed to explore their capabilities and limitations.
In the context of IDE models, some significant strides have been made in constraining model parameters using the galaxy power spectrum. A notable study by  Nunes et al. (2022a) provided an in-depth analysis of the coupling parameter $\xi$, which represents the interaction strength between DE and DM. By combining the full power spectrum from the BOSS DR12 galaxy sample with Planck CMB data, the authors established a lower bound of $\xi>-0.12$ at the 95% CL. Despite this progress, there remains a paucity of studies that employ non-linear modeling within the IDE framework using real observational data, highlighting a critical area for future research.
Given the imperative to anticipate model behavior on non-linear scales, it becomes essential to employ phenomenological models for the interpretation of current and forthcoming data derived from galaxy clustering and cosmic shear. Exploring this approach is particularly important as it helps bridge the gap between theoretical predictions and observational data. By doing so, we can enhance our understanding of the perturbation dynamics within these complex systems and improve our comprehension of the observational data. In this work, our main aim is to build a robust non-linear $P(k)$ model in the presence of a DE-DM interaction. To this end, we carry out a series of N-body gravity simulations and then we develop these perspectives in two ways. First, we capture the ratio $R(k)=P_{\rm NL}(k)/P_{\rm linear}(k)$ up to non-linear scales, and we find reliable analytical modeling for the function $R(k)$ up to scales $k\leq 1$ h/Mpc. In a second modeling context, we modify and propose a new $P(k)$ Halo model, which can fit the simulation data. Finally, we applied our models in a forecast analysis from the perspective of the Euclid survey. We demonstrate that an Euclid-like survey will possess strong sensitivity to constrain DM-DE coupling, rendering it a crucial probe for this particular class of cosmological models.
@@ II Background expansion and linear perturbations
The phenomenological possibility of a dark sector interaction has been extensively investigated over the past two decades. Pioneering work in this area was presented in Amendola (2000); Farrar and Peebles (2004); Zimdahl et al. (2001); Baldi (2010); Nusser et al. (2005); Amendola et al. (2008), and it has since been extended to observational tests Kumar and Nunes (2016); Murgia et al. (2016); Kumar and Nunes (2017); Di Valentino et al. (2017); Yang et al. (2019a, 2018, b); Kumar et al. (2019); Pan et al. (2019a, b); Di Valentino et al. (2020a, b); Di Valentino and Mena (2020); Di Valentino et al. (2021c); Gómez-Valent et al. (2020); Yang et al. (2020a); Lucca and Hooper (2020a); Martinelli et al. (2019); Yang et al. (2020b); Yao and Meng (2020); Pan et al. (2020a); Di Valentino (2020); Yao and Meng (2021); Amirhashchi et al. (2022); Yang et al. (2021a); Gao et al. (2021); Lucca (2021a); de Cruz Perez and Sola Peracaula (2024); Kumar (2021); Yang et al. (2021b); Lucca (2021b); Halder and Pandey (2021); Bernui et al. (2023); Poulin et al. (2023); Lucca (2021c); Zhai et al. (2023); Gariazzo et al. (2022); Valentino et al. (2020); Mawas et al. (2021); Yang et al. (2021c); Lucca and Hooper (2020b); Pan et al. (2020b); von Marttens et al. (2020); van der Westhuizen and Abebe (2024); Goh et al. (2023); Gómez-Valent et al. (2022); Escamilla et al. (2023); Forconi et al. (2023); Hoerning et al. (2023); Wang et al. (2024a); Benisty et al. (2024) and N-body simulations Macciò et al. (2004); Baldi et al. (2010). For a comprehensive bibliographic review, see Wang et al. (2024b, 2016). In this section, we will focus on recent developments essential to establishing the theoretical framework central to this work.
Let us start by examining the fundamental characteristics of IDE models. Throughout our discussion, we will operate on the assumption of a spatially flat Friedmann-Lemaître-Robertson-Walker (FLRW) metric.
We can introduce a phenomenological parameterization for the DM-DE interaction within the conservation equations such that the individual stress-energy tensors are no longer conserved, although their sum remains so. [^1] See Nunes et al. (2022b) for a more detailed description and Wang et al. (2024b, 2016) for a general review.
Typically, it is assumed that the covariant derivatives of the DM and DE stress-energy tensors evolve according to:
@@
The introduction of the DM-DE coupling further alters the evolution of perturbations. Working in the Newtonian gauge, the coupled system of linear Einstein-Boltzmann equations governing the evolution of dark matter density perturbations, $\delta_{c}$, and velocity divergences, $\theta_{c}$, is expressed as follows Valiviita et al. (2008); Honorez et al. (2010); Gavela et al. (2010):
@@
It is worth noting that the Euler equation remains unaltered compared to $\Lambda$CDM, indicating the absence of a fifth force in this model. Additionally, we will not consider the potential presence of clustering DE in this study. As argued in Dakin et al. (2019), DE perturbations may only have discernible effects on very large scales, necessitating a relativistic treatment of simulations. In the regime of Newtonian N-body simulations that we are focused on modeling, DE clustering does not significantly impact the matter power spectrum Hassani et al. (2019). Moreover, recent analyses Dinda and Banerjee (2023) have demonstrated that observations favor a homogeneous DE component over a clustering one. However, the presence of a homogeneous DE component influences simulations in various ways, including background evolution, growth function calculation, and initial conditions. To provide a more precise quantification within the IDE context, Figure 1 illustrates the relative difference in $P(k,z=0)$ with and without DE clustering effects. In constructing the figure, we adopted IDE-01 and IDE-02 baseline values, which will be defined in the subsequent section. It is evident that on large scales, the observed effect is minimal, amounting to less than 1%. As for scales where non-linear effects are expected to dominate (i.e., $k>0.2h/\text{Mpc}$), the relative difference is approximately 1%. Consequently, we can infer that the perturbations due to DE can be considered negligible. The small oscillations observed in the curves at large scales, often perceived as noise, are attributed to the intrinsic oscillatory nature of $P(k)$ when numerically calculating the ratio between the desired quantities. These fluctuations are negligible from a practical standpoint, as they result in a relative difference of approximately 0.001% or less. The results were obtained using a customized version of the CLASS code Blas et al. (2011), which integrates the IDE model elaborated upon in this study.
@@
Figure 2: The growth factor is depicted as a function of redshift. The dashdot black curve represents the standard model, $\Lambda$CDM, whereas the solid curves correspond to the interacting model utilized, showcasing variations of the $\xi$ parameter.
@@ III.1 Methodology
The cosmic structure formation in a given theory is most accurately modeled with N-body simulations, where the particles in the simulation are incrementally displaced from their initial positions according to the Poisson equation, the expansion rate of the universe and initial conditions. In this work, we will use the pmwd code Li et al. (2022), a differentiable cosmological particle-mesh N-body library. On the other hand, codes based on the Particle-Mesh method, like the pmwd code, may not be as robust as well-established routines, such as those implemented in Gadget Springel (2005); Springel et al. (2021), at very small scales. However, when conducting two separate analyses with a finer grid resolution within the same volume, we observed no significant differences in the predicted non-linear power spectrum, $P_{\rm NL}(k)$, within the scales of interest for this work. It is important to note that all our simulations for extracting $P_{\rm NL}(k)$ exclusively involve dark matter particles.
@@
To test our framework for the IDE model in N-body simulations introduced above, we shall consider three scenarios: the $\Lambda$CDM model from Planck-CMB best-fit values, and, based on the recent LSS analysis Nunes et al. (2022a) in the IDE context, two different input baselines for IDE. Therefore, we will consider the following common fiducial values: $2.110^{9}A_{s}=$ 2.1, $n_{s}=0.96$, and $\Omega_{b}=0.05$, while we will choose for:
@@ III.2 Discussion
In the following discussion, the superscript $s$ indicates quantities evaluated differently from N-body simulation data. Figure 4 illustrates the ratio between $R_{\text{IDE}}^{s}(k)$ and $R_{\text{IDE}}(k)$, parameterized by equation (31). Our analysis, through chi-square minimization, reveals that the parameters $A\in[23.23,24.24]$, $B\in[3.77,4.74]$, $C\in[-0.03,-0.04]$, and $D\in[-9.30,-16.35]$ can effectively fit the non-linear spectra within a $<$5% relative difference for the values within the range adopted in IDE-01 and IDE-02 samples. It is noteworthy that any observed noise in Figure 4 on very large scales is solely attributable to the box size of our simulations. Thus, for the regime where non-linear effects are predominant, i.e., $k>0.2$ h/Mpc, we note that our fit is very precise with accuracy $<$1%.
@@ IV The Halo Model for IDE
The leading order moment of the cosmological matter distribution is the non-linear power spectrum, $P_{\rm NL}(k)$. The development and results presented in the previous section offer a parametric function that relates the predictions $P_{\rm NL}(k)$ and $P_{\rm linear}$. While this approach is commonly used in the literature, we also focus on a second argument: the development of a new halo model that can incorporate the model of interest in this work.
To do this, we used the structure based on the HMCode Mead et al. (2016) (see Asgari et al. (2023) for a review of halo models). By incorporating simulation-calibrated quantities, such as the halo mass function and dark density profiles adjusted for baryonic feedback, the halo model offers a robust integration between theoretical predictions and observational data. This approach is especially effective in modeling the nonlinear power spectrum, $P_{\rm NL}(k)$, and allows the model to be reused for other analyses without the need for additional simulations, which are expensive in terms of cost and computation.
@@
In the above equation, the power spectrum is formulated as an integral over halo masses, $M$, with the differential halo mass function denoted as $dM$, determined by the halo mass function, $F(M)$, which predicts the abundance of dark matter halos of different masses. Additionally, this integral incorporates the term $W(k,M)$, representing the normalized Fourier transform of the halo density profile Asgari et al. (2023). On larger spatial scales, the distribution of haloes exhibits non-random patterns, and the displacements between them necessitate the incorporation of a two-halo term into the power spectrum analysis. This term accounts for the interactions between haloes and is akin to the linear-theory power spectrum for the matter distribution, $\Delta_{2\mathrm{H}}^{2}(k)=\Delta_{\mathrm{lin}}^{2}(k)$.
@@
Our analysis indicates that $\beta_{1}\in[3.4,3.6]$, $\beta_{2}\in[0.15,0.28]$, $\beta_{3}\in[0.15,1.31]$, $\beta_{4}\in[3.5,5.1]$ and $\beta_{5}\in[1.09,1.11]$ can fit with good accuracy the entire parameter-space range between the IDE-01 and IDE-02 model baselines. The determination of the fitting values involved an extensive study of the HMCode already implemented in CLASS, where parameter variations were analyzed to assess the impact of the IDE cosmology on the key functions of the adopted halo model. The creation of these parameters played a crucial role in our model’s ability to accurately describe the power spectra of matter on non-linear scales. Figure 5 illustrates the ratio between the modified halo model and simulation data for two scenarios (IDE-01 and IDE-02). Notably, modeling accuracy within a 5% margin extends to values of $k\approx 1$. A relative difference of less than 5% compared to simulations is a promising outcome, suggesting that our model may be robust enough to capture the subtleties of matter structures even in different settings, as tested with IDE-01 and IDE-02. It is important to note that the Euclid survey aims to measure power spectra with 1% accuracy at linear scales. Our current calibration achieves an adjustment within 5% for scales up to approximately 1 h/Mpc. Given the complexity of the phenomenological model under consideration in this work, and considering that our proposal is pioneering—specifically, in developing a new halo model for the IDE scenario—we plan to offer further refinements in a future publication, incorporating a significantly larger number of simulations and enhanced precision.
@@ V Forecast analysis from the Euclid Survey perspective
Parameter $\mathbf{\sigma(\text{Euclid})}$
$\mathbf{k_{max}=0.1}$ $\mathbf{k_{max}=0.5}$ $\mathbf{k_{max}=1}$
$\Omega_{\rm cdm}$ 0.037 0.018 0.015
$n_{s}$ 0.019 0.0065 0.0055
$h$ 0.0169 0.0085 0.0066
$\xi$ 0.0136 0.0131 0.0110
$\sigma_{\rm NL}$ 0.137 0.072 0.059
$b_{0}$ 0.0108 0.0039 0.0032
$b_{1}$ 0.0143 0.0066 0.0058
@@
In what follows in our main results, we use a modified version of the CLASS$+$MontePython code Blas et al. (2011); Audren et al. (2013); Brinckmann and Lesgourgues (2018), adapting the likelihoods euclid_pk as well as incorporating all background dynamics, linear perturbations, and the modified halo models in CLASS, as discussed in the previous sections.
In our forecast analyses, we considered three cases by applying different cut-off scales: the linear regime ($k_{\rm max}=0.1h/\text{Mpc}$), the mildly non-linear ($k_{\rm max}=0.5h/\text{Mpc}$), and the non-linear ($k_{\rm max}=1h/\text{Mpc}$). Our model exhibits good predictive capability across these scales, successfully fitting the simulations. Furthermore, we incorporate baryon feedback using a six-parameter physical model, which includes AGN-driven gas ejection and star formation processes, as implemented in HMCode Mead et al. (2020). This model, validated against hydrodynamic simulations, accurately captures feedback effects in the nonlinear power spectrum with sub-percent precision, even on very small scales. For a recent discussion on the significance of baryonic effects in the power spectrum, see also Parimbelli et al. (2019); Chisari et al. (2019); Terasawa et al. (2024); Xu et al. (2023); Salcido et al. (2023).
@@
Of particular significance is the examination of sensitivity towards the coupling parameter $\xi$. According to the analysis of Euclid forecasts, the coupling parameter can be constrained with a precision of $\sigma(\xi)=0.013$ (0.011) considering only linear (non-linear) scales, respectively. This represents a substantial improvement by an order of magnitude compared to any other current observational tests documented in the literature. For example, in the study by Nunes et al. (2022a), which employed the full-form galaxy power spectrum of the BOSS DR12 sample in conjunction with CMB data from Planck, a constraint of $\xi>-0.12$ was established. Consequently, the forthcoming data from Euclid has the potential to enhance these constraints by up to two orders of magnitude. Therefore, the accuracy of the constraints generated by the Euclid mission sensitivity will be crucial for the IDE models. The joint accuracy with $\sigma(\Omega_{\rm cdm})$ and $\sigma(h)$, two parameters highly sensitive to IDE dynamics, will also be constrained with remarkable precision. Consequently, Euclid data alone will possess the capability to constrain the dynamics of IDE models, encompassing both background and perturbative levels, with unprecedented accuracy.
@@
There are other noteworthy points to consider. In forecast analyses from the Euclid perspective, a persistent and significant negative correlation between the coupling parameter $\xi$ and the Hubble constant $h$ is observed. This correlation has been pivotal in supporting the argument that this class of models might resolve the $H_{0}$ tension Zhai et al. (2023); Bernui et al. (2023); Gariazzo et al. (2022). Interestingly, the analysis also suggests that $\xi$ may exhibit a notable new correlation with the galaxy bias parameters $b_{0}$ and $b_{1}$ (see Figure 6). Moreover, the observed correlation in the $\xi$-$h$ plane within the IDE framework implies that the galaxy bias parameters are likely to correlate with $h$ as well. This insight unveils a previously unrecognized cosmological characteristic of these models, which will be crucial to understand in future analyses involving real observational data.